# Heterogeneous network estimation for single-cell transcriptomic data via a joint regularized deep neural network

### v1

@@ 4.2 JRDNN-KM constructs heterogeneous networks with common edges
The inferred networks with JRDNN-KM are shown in Figure 5 for the LUAD dataset and Supplementary Figures S6-S9 for the rest four datasets. The identified numbers of edges for different subgroups are reported in Supplementary Figure S10-S14, where 70, 88, 76, 2, and 0 common edges are shared by all subgroups for the LUAD, PBMC, mESCs, and mouse liver and uterus datasets, respectively. The numbers of common edges for the mouse liver and uterus datasets are smaller compared to the other three datasets. This is reasonable, as the corresponding cell subgroups have relatively larger diversities. We take a closer look at the Kupffer and T cell subgroups for the mouse liver dataset, which both participate in the regulation process of the immune system, and identify 71 common edges. For the mouse uterus dataset, we also examine the two immune regulation-related cell subgroups, Macrophage and NK cells, and identifies 26 common edges. In addition, 28 common edges are shared by the Endothelial and Muscle cells, which both participate in maintaining the normal physiological functions of the body. We also conduct analysis with the alternative methods, where the true cell subgroup memberships are adopted for JGNsc, JSEM, SpQN, Normalisr, GENIE3, locCSN, and CS-CORE, and present the comparison results in Supplementary Figure S10-S14. It can be seen that the nine methods identify a moderate number of overlapping edges, where the overlapping ratio of the identified edges between the proposed method and others lies in [52.7%,82.8%], and JRDNN-KM can effectively exploit the common information across different subgroups.
@@ 4.3 Investigation of hub genes for the LUAD data
We further take the LUAD data from humans as an example and investigate the hub genes in the inferred networks with JRDNN-KM. The hub genes are important topological features of networks that usually have functional relevance. In this study, we consider the top ten genes with the largest degrees in the networks (shown in Supplementary Table S4 and highlighted with a bigger point in Figure 5 and Supplementary Figures S6-S9). It can be seen that JRDNN-KM identifies two common hub genes (RPR11 and ASPM) shared by all three cell lines (H1975, H2228, and HCC827) and nine common hub genes shared by two of the three cell lines. The similarity across these three cell lines has been universally recognized in the literature. In addition, two specific hub genes (FTH1 and HSP90AB1), one specific gene (ARL6IP1), and three specific genes (MT-ND4, AKR1C3, and TPM4) are identified for subgroup 1 (H1975), subgroup 2 (H2228), and subgroup 3 (HCC827), respectively.
@@
The top ten hub genes identified with the alternatives and the comparison results among different methods are also reported in Supplementary Table S4. It is observed that JRDNN-KM detects a few well-known marker genes that are found by the majority of approaches and also reveals some promising novel findings. Specifically, the two common hub genes, RPL11 and ASPM, shared by all three cell lines, are also recognized by most of the competing methods. Here, ASPM has emerged as a critical player in cancer biology, specifically influencing cancer aggressiveness and stemness. ASPM’s overexpression in various malignancies has been demonstrated to be strongly associated with poor patient outcomes, underscoring its role in oncogenic processes and its impact on cancer cell behavior and tumor evolution (Tsai et al., 2023). In addition, RPL11 is crucial in cancer progression through its interaction with the MDM2-P53 pathway. In particular, RPL11 acts as a tumor suppressor by stabilizing P53, particularly when PICT1 is deficient with the inhibition of MDM2’s activity, which correlates with slower tumor growth and potentially improved patient outcomes (Sasaki et al., 2011).
@@
We further examine the subgroup-specific hub genes, where JRDNN-KM demonstrates its unique capability by recognizing hub genes that most competing methods missed. For instance, in the H1975 cell line, the FTL gene and in the HCC827 cell line, the AKR1C3 gene are only identified by JRDNN-KM and locCSN. These methods, both employing non-linear approaches, effectively highlight the non-linear regulatory relationships among certain genes in LUAD, which might be overlooked before. Additionally, the CFL1 gene in the H2228 cell line is recognized only by JRDNN-KM. Biological implications of these genes have been well recognized in the literature. Specifically, the FTL gene plays a pivotal role in regulating ferroptosis by controlling iron storage and reducing oxidative stress under the governance of the NRF2 pathway. This regulation helps LUAD cells evade ferroptosis, thereby contributing to potential treatment resistance (Ding et al., 2024). AKR1C3 is consistently overexpressed in tumor tissues, suggesting its significant association with the disease and indicating its involvement in tumor progression and resistance to erlotinib therapy in LUAD (Cho et al., 2024). In addition, the CFL1 gene and its functional gene network have been suggested as prognostic biomarkers for lung adenocarcinoma, which can also guide chemotherapeutic interventions (Castro et al., 2010).
@@ 4.4 Community detection and GO enrichment analysis for the LUAD data
We continue to conduct community (module) detection on the referred networks using the Louvain algorithm and show two communities with the largest sizes in Figure 5 and Supplementary Figures S6-S9. Both homogeneous and heterogeneous edges are observed in these communities. Similar to the analysis on the hub genes, a deeper examination based on the gene ontology (GO) enrichment analysis is conducted for the LUAD data, where the top five GO terms with the smallest P-values for each subgroup are presented in Figure 6. The results suggest that the communities share some significantly enriched GO terms, indicating functional and biological connections among the related genes. The community detection analysis is also conducted for the competing methods, and the top ten significant GO terms involved in the two largest communities are reported in Supplementary Tables S5-S7. It can be seen that some GO terms are shared by all the nine methods. Compared to the H1975 and H2228 cell lines, JRDNN-KM has more unique findings in the HCC827 cell line.
@@
Specifically, as shown in Figure 6, “cytoplasmic translation” is among the top five significant GO terms in all three subgroups. This biological process has been found to be crucial as it facilitates the synthesis of proteins necessary for cancer cell growth and proliferation and is notably involved in the progression of LUAD. In addition, the p53 class mediator-related GO terms are significantly enriched in both H2228 and HCC827 cell lines. Although not featured among the top five significant terms, these terms are also significant in the H1975 cell line, specifically the “regulation of signal transduction by p53 class mediator” with a p-value of $4.736\times 10^{-4}$, and “positive regulation of signal transduction by p53 class mediator” with a p-value of $7.894\times 10^{-4}$. In LUAD, p53 has been shown to suppress tumor development by promoting the differentiation of alveolar type 1 (AT1) cells from transitional states during alveolar repair, effectively governing cell state to prevent the persistence of cancerous cell types and maintain lung tissue integrity (Kaiser et al., 2023). These two terms are also detected by all competing methods with significant p-values.
@@
Similar to the analysis of hub genes, JRDNN-KM identifies some unique GO terms, such as “regulation of translation” found only by JRDNN-KM and locCSN in the H1975 cell line and “regulation of protein modification” discovered only by JRDNN-KM in the HCC827 cell line, suggesting that JRDNN-KM reveals relevant novel biological processes. Recent research shows that in the H1975 cell line, significant alterations in translation regulators were observed, particularly in rociletinib-resistant cells. These changes may affect the synthesis of key proteins, potentially influencing tumor progression and response to therapy (Zhang et al., 2021). And in HCC827 cells, inhibiting N-linked glycosylation, a key protein modification, disrupts EGFR function, reducing proliferation and inducing senescence (Lopez-Sambrooks et al., 2016). These unique findings of JRDNN-KM correlate well with practical biological explanations, providing valuable insights for research and clinical applications.
Figure 6: GO enrichment analysis results of two representative communities identified with JRDNN for the LUAD data. (A) Subgroup 1, (B) Subgroup 2, and (C) Subgroup 3.
@@ 5 Discussion
In this study, we have proposed JRDNN-KM, a systemic statistical framework for estimating heterogeneous networks and cell subgroups from single-cell transcriptomic data. JRDNN-KM has leveraged the robust fitting capabilities of neural networks to approximate the complex nonlinear relationships between genes and incorporated the inferred networks into the K-means clustering to accommodate both the mean and covariance heterogeneous among cell subgroups. JRDNN-KM has directly addressed the challenging characteristics of single-cell transcriptomic data, including the nonlinearity, zero-inflation, unknown cellular heterogeneity, as well as a certain degree of homogeneity among cell subgroups. Under extensive simulation studies, JRDNN-KM has demonstrated substantial improvements over recent network estimation models, particularly in handling complex nonlinear relationships among genes that are not adequately captured by models such as GGMs.
Our analyses of real single-cell transcriptomic data from various sources have further validated the utility of JRDNN-KM. Besides the accurate cell subgroup identification performance, the networks identified by JRDNN-KM have been consistently corroborated by biological research. On one hand, the biologically implicated common hub genes and significant GO terms shared by different cell subgroups highlight JRDNN-KM’s potential in investigating the homogeneous dependencies and interactions among genes in the same tissue. On the other hand, the heterogeneous topological properties of the inferred networks providing deeper insights into the subgroup specific-cellular mechanisms underlying different biological states and conditions. These two different types of characteristics are crucial for practical applications in biomedical research and therapy development.
Our study, while groundbreaking, acknowledges certain limitations. For instance, our proposed loss function is based on squared loss, which may lack robustness against outliers and noise. Future work could explore optimizing this aspect to enhance the model’s stability and performance under varying conditions. Additionally, our assumptions regarding zero inflation are based on varying zero-inflation rates across different genes. However, emerging research indicates that technical dropout may be correlated with expression levels, suggesting a need for a more nuanced approach to account for these dependencies in our model. Due to computational constraints, our study necessitates preprocessing the data before applying our JRDNN-KM method. Future work will concentrate on refining the model to enhance its capacity to handle even larger datasets, thereby broadening its applicability across various areas of genomics.
@@
Funding
This research was supported by the National Natural Science Foundation of China (12071273); Shanghai Rising-Star Program (22QA1403500); Shanghai Science and Technology Development Funds (23JC1402100); Shanghai Research Center for Data Science and Decision Technology; National Institutes of Health (CA204120 and CA121974); and National Science Foundation (2209685).
@@
Supplementary Materials
Additional supporting information may be found online in the supplementary materials section at the end of the article.
Supplement.pdf
Supplementary Materials for “Heterogeneous network estimation for single-cell transcriptomic data via a joint regularized deep neural network”, including the details of the proposed algorithms and additional settings and results of simulation studies and real data analysis.
JRDNN-KM.zip
The package JRDNN-KM that implements the proposed approach, which is available at https://github.com/mengyunwu2020/JRDNN-KM.

### v2

@@ 4.3 JRDNN-KM constructs heterogeneous networks with common edges
The inferred networks with JRDNN-KM are shown in Figure 4 for the LUAD dataset and Supplementary Figures S7-S10 for the rest four datasets. The identified numbers of edges for different subgroups are reported in Supplementary Figures S11-S15, where 70, 88, 76, 2, and 0 common edges are shared by all subgroups for the LUAD, PBMC, mESCs, and mouse liver and uterus datasets, respectively. The numbers of common edges for the mouse liver and uterus datasets are smaller compared to the other three datasets. This is reasonable, as the corresponding cell subgroups have relatively larger diversities. We take a closer look at the Kupffer and T cell subgroups for the mouse liver dataset, which both participate in the regulation process of the immune system, and identify 71 common edges. For the mouse uterus dataset, we also examine the two immune regulation-related cell subgroups, Macrophage and NK cells, and identifies 26 common edges. In addition, 28 common edges are shared by the Endothelial and Muscle cells, which both participate in maintaining the normal physiological functions of the body. Comparative results across methods are provided in Supplementary Figures S11-S15. It can be seen that the nine methods identify a moderate number of overlapping edges, where the overlapping ratio of the identified edges between the proposed method and others lies in [52.7%,82.8%], and JRDNN-KM can effectively exploit the common information across different subgroups.
@@ 4.4 Investigation of hub genes for the LUAD data
We further take the LUAD data from humans as an example and investigate the hub genes in the inferred networks with JRDNN-KM. The hub genes are important topological features of networks that usually have functional relevance. In this study, we consider the top ten genes with the largest degrees in the networks (shown in Supplementary Table S4 and highlighted with a bigger point in Figure 4 and Supplementary Figures S7-S10). It can be seen that JRDNN-KM identifies two common hub genes (RPR11 and ASPM) shared by all three cell lines (H1975, H2228, and HCC827) and nine common hub genes shared by two of the three cell lines. The similarity across these three cell lines has been universally recognized in the literature. In addition, two specific hub genes (FTH1 and HSP90AB1), one specific gene (ARL6IP1), and three specific genes (MT-ND4, AKR1C3, and TPM4) are identified for subgroup 1 (H1975), subgroup 2 (H2228), and subgroup 3 (HCC827), respectively.
@@
The top ten hub genes identified with the alternatives and the comparison results among different methods are also reported in Supplementary Table S4. It is observed that JRDNN-KM detects a few well-known marker genes that are found by the majority of approaches and also reveals some promising novel findings. Specifically, the two common hub genes, RPL11 and ASPM, shared by all three cell lines, are also recognized by most of the competing methods. Here, ASPM has emerged as a critical player in cancer biology, specifically influencing cancer aggressiveness and stemness. ASPM’s overexpression in various malignancies has been demonstrated to be strongly associated with poor patient outcomes, underscoring its role in oncogenic processes and its impact on cancer cell behavior and tumor evolution. In addition, RPL11 is crucial in cancer progression through its interaction with the MDM2-P53 pathway. In particular, RPL11 acts as a tumor suppressor by stabilizing P53, particularly when PICT1 is deficient with the inhibition of MDM2’s activity, which correlates with slower tumor growth and potentially improved patient outcomes.
@@
We further examine the subgroup-specific hub genes, where JRDNN-KM demonstrates its unique capability by recognizing hub genes that most competing methods missed. For instance, in the H1975 cell line, the FTL gene and in the HCC827 cell line, the AKR1C3 gene are only identified by JRDNN-KM and locCSN. These methods, both employing non-linear approaches, effectively highlight the non-linear regulatory relationships among certain genes in LUAD, which might be overlooked before. Additionally, the CFL1 gene in the H2228 cell line is recognized only by JRDNN-KM. Biological implications of these genes have been well recognized in the literature. Specifically, the FTL gene plays a pivotal role in regulating ferroptosis by controlling iron storage and reducing oxidative stress under the governance of the NRF2 pathway. This regulation helps LUAD cells evade ferroptosis, thereby contributing to potential treatment resistance. AKR1C3 is consistently overexpressed in tumor tissues, suggesting its significant association with the disease and indicating its involvement in tumor progression and resistance to erlotinib therapy in LUAD. In addition, the CFL1 gene and its functional gene network have been suggested as prognostic biomarkers for lung adenocarcinoma, which can also guide chemotherapeutic interventions. References supporting the discussions on biological functionalities are provided in Supplementary Section S5.
@@ 4.5 Community detection and GO enrichment analysis for the LUAD data
We continue to conduct community (module) detection on the referred networks using the Louvain algorithm and show two communities with the largest sizes in Figure 4 and Supplementary Figures S7-S10. Both homogeneous and heterogeneous edges are observed in these communities. Similar to the analysis on the hub genes, a deeper examination based on the gene ontology (GO) enrichment analysis is conducted for the LUAD data, where the top five GO terms with the smallest P-values for each subgroup are presented in Supplementary Figure S16. The results suggest that the communities share some significantly enriched GO terms, indicating functional and biological connections among the related genes. The community detection analysis is also conducted for the competing methods, and the top ten significant GO terms involved in the two largest communities are reported in Supplementary Tables S5-S7. It can be seen that some GO terms are shared by all the nine methods. Compared to the H1975 and H2228 cell lines, JRDNN-KM has more unique findings in the HCC827 cell line.
@@
Specifically, as shown in Supplementary Figure S16, “cytoplasmic translation” is among the top five significant GO terms in all three subgroups. This biological process has been found to be crucial as it facilitates the synthesis of proteins necessary for cancer cell growth and proliferation and is notably involved in the progression of LUAD. In addition, the p53 class mediator-related GO terms are significantly enriched in both H2228 and HCC827 cell lines. Although not featured among the top five significant terms, these terms are also significant in the H1975 cell line, specifically the “regulation of signal transduction by p53 class mediator” with a p-value of $4.736\times 10^{-4}$, and “positive regulation of signal transduction by p53 class mediator” with a p-value of $7.894\times 10^{-4}$. In LUAD, p53 has been shown to suppress tumor development by promoting the differentiation of alveolar type 1 (AT1) cells from transitional states during alveolar repair, effectively governing cell state to prevent the persistence of cancerous cell types and maintain lung tissue integrity. These two terms are also detected by all competing methods with significant p-values.
@@
Similar to the analysis of hub genes, JRDNN-KM identifies some unique GO terms, such as “regulation of translation” found only by JRDNN-KM and locCSN in the H1975 cell line and “regulation of protein modification” discovered only by JRDNN-KM in the HCC827 cell line, suggesting that JRDNN-KM reveals relevant novel biological processes. Recent research shows that in the H1975 cell line, significant alterations in translation regulators were observed, particularly in rociletinib-resistant cells. These changes may affect the synthesis of key proteins, potentially influencing tumor progression and response to therapy. And in HCC827 cells, inhibiting N-linked glycosylation, a key protein modification, disrupts EGFR function, reducing proliferation and inducing senescence. These unique findings of JRDNN-KM correlate well with practical biological explanations, providing valuable insights for research and clinical applications (References supporting these discussions are provided in Supplementary Section S5).
4.6 Sensitivity and Subsampling Analysis
[…]
Moreover, to evaluate the robustness of our model, we perform a subsampling analysis across the five real datasets. Specifically, we randomly subsample 80% of the data from each dataset and repeat this process 50 times to create independent replicates. As shown in Supplementary Figures S18-S21, our proposed method consistently maintains its performance advantage and exhibits high stability across the subsampled replicates.
5 Application-Based Simulation
In this section, we provide application-based simulations to evaluate our method using three benchmark datasets: the Gonadal Sex Determination (GSD) dataset (Pratapa et al., 2020), scMultiSim-T3 dataset (Li et al., 2025), and SERGIO-DS1 dataset (Dibaeinia and Sinha, 2020). These datasets were derived from real biological networks and incorporate practical biological processes, offering a challenging and biologically realistic benchmark for testing network inference algorithms. Specifically, GSD dataset is based on a Boolean network model of gonadal sex determination, which captures the bipotential differentiation of gonads into either male (Sertoli cells) or female (Granulosa cells) fates. scMultiSim-T3 dataset is generated using scMultiSim, a multi-modality single-cell data simulator that integrates biological factors including gene regulatory networks (GRNs) and cell-cell interactions (CCIs). SERGIO-DS1 dataset is generated by the single-cell gene expression simulator SERGIO, which is constructed based on a GRN derived from Ecoli. More detailed descriptions of these datasets are provided in Supplementary Section S6.
Similar to real data analysis, we report the summarized ARI, F1, Recall, and Precision values in Figure 5. JRDNN-KM demonstrates a clear and consistent advantage over competing methods. It achieves the highest ARI for cell subgroup identification and also excels in gene network inference. The performance of all methods generally declines from GSD to scMultiSim-T3 to SERGIO-DSI, indicating that the SERGIO-DSI dataset presents the greatest challenge. In this most difficult scenario, JRDNN-KM maintains its relative effectiveness. While its ARI (0.834) and F1 score (0.481) on SERGIO-DSI are lower than its results on the other datasets, it maintains a similar performance margin over the alternatives. For instance, its lead in F1 score over the second-best method, BLGGM, remains approximately 0.027, 0.027, and 0.024 on GSD, scMultiSim-T3, and SERGIO-DSI, respectively. The observed consistent advantage confirms that the joint learning of subgroups and networks in JRDNN-KM remains effective even in complex and noisy data environments.
Figure 5: Comparative performance of different methods on three application-based simulation datasets; (A)-(D): ARI, F1, Recall, and Precision values.
We also perform a comprehensive set of simulation studies (Supplementary Section S7) to assess method performance under diverse conditions, including variations in network structure, nonlinearity, common information, dropouts, and sample distributions. JRDNN-KM again proves superior, outperforming other methods in accurately identifying cell subgroups and estimating gene networks.
@@ 6 Discussion
Estimating heterogeneous gene networks remains a critically important research focus in biomedical science (Yang et al., 2025). In this study, we have proposed JRDNN-KM, a systemic statistical framework for estimating heterogeneous networks and cell subgroups from single-cell transcriptomic data. JRDNN-KM leverages the robust fitting capabilities of neural networks to approximate complex nonlinear gene-gene relationships and integrates the inferred networks into K-means clustering, thereby accounting for heterogeneity in both gene expression levels and gene network structures across cell subgroups. JRDNN-KM has directly addressed the challenging characteristics of single-cell transcriptomic data, including the nonlinearity, zero-inflation, unknown cellular heterogeneity, as well as a certain degree of homogeneity among cell subgroups.
Our analyses of real single-cell transcriptomic data from various sources have validated the utility of JRDNN-KM. Besides the accurate cell subgroup identification performance, the networks identified by JRDNN-KM have been consistently corroborated by biological research. We have analyzed LUAD data as an illustrative case, performing downstream hub gene and community detection analyses based on the estimated networks, which are two pivotal network topological features. Subsequent GO term enrichment analysis has been applied to the two largest communities. As shown in Supplementary Tables S4-S7, our method has uniquely revealed functionally coherent, literature-validated targets such as FTL in ferroptosis regulation and GO:1903320 in regulation of protein modification, which were overlooked by alternative approaches. They offer new insights into homogeneous gene dependencies within tissues alongside subgroup-specific cellular mechanisms governing distinct biological states. As candidate biomarkers, these hub genes and GO term-associated biological functions potentially facilitate target discovery and therapeutic hypothesis generation. Extensive simulation studies have further demonstrated the substantial improvements of JRDNN-KM over recent network estimation models, particularly in handling complex nonlinear relationships among genes that are not adequately captured by models such as GGMs.
In our study, the proposed loss function is based on squared loss, which may lack robustness against outliers and noise. Future work could explore optimizing this aspect to enhance the model’s stability and performance under varying conditions. Our analysis utilizes zero-inflated models in accordance with established methods. Alternative approaches employing Poisson or negative binomial distributions, particularly when integrated with deep neural network architectures, represent a compelling avenue for future methodological innovation. Additionally, our assumptions regarding zero inflation are based on varying zero-inflation rates across different genes. However, emerging research indicates that technical dropout may be correlated with expression levels, suggesting a need for a more nuanced approach to account for these dependencies in our model. In heterogeneous analysis, the indicator matrix $\boldsymbol{\Theta}_{k}$ is used within the Mahalanobis distance framework. Our study is designed to detect nonlinear relationships between variables rather than quantify their precise strength. Consistent with most literature on DNNs, precisely estimating effect magnitudes remains challenging due to the inherent “black-box” nature of deep architectures. Future work could further explore quantifying the strength of nonlinear relationships. Our method focuses on point estimates of network structure, consistent with standard practice in the literature, including graphical lasso (Zhang and Zou, 2014; Dong et al., 2023), neighborhood selection (Ma and Michailidis, 2016), and deep learning approaches (Shu et al., 2021). While uncertainty quantification would be valuable, it remains challenging for high-dimensional nonlinear frameworks like ours and is deferred to future study. To enhance computational efficiency, we adopt the shared parameter framework, thereby avoiding estimation of nuisance parameters $\sigma_{kj}^{2}$. This approach has demonstrated satisfactory performance in both simulation studies and real data analyses. Future work could explore estimating $\sigma_{kj}^{2}$ to enable heterogeneous penalty parameterization.
In real data analysis, for network validation, we have employed four established reference networks. Although ChIP-seq data provide valuable evidence of direct TF-DNA binding, its utility in our 100-gene context is limited by the sparse representation of TF: too few transcription factors with validated targets were included to form meaningful regulatory connections. Here, we have focused on 100 HV genes, a common practice in existing network estimation studies. Given that we need to consider relationships among $p$ variables (i.e., $p\times(p-1)$ potential edges), an excessively large $p$ would lead to high computational complexity. This has also been a challenging issue in traditional GGMs. In future work, we will focus on optimizing the algorithm to improve model scalability.
@@
Supplementary Materials
In supplementary materials, the pdf file contains supplementary sections, supplementary figures, and supplementary tables referred in the paper. The zipped file includes the code and datasets used in the simulation and real application and provides detailed instructions for reproducing results. The code that implements the proposed approach is also available at https://github.com/mengyunwu2020/JRDNN-KM.
Acknowledgments
The authors thank the editors and reviewers for their invaluable feedback and insightful suggestions, which have significantly improved this paper.
@@ Acknowledgments
Data Availability Statement
The data that support the findings of this study are openly available in https://github.com/mengyunwu2020/JRDNN-KM.
Funding
This research was supported by the National Natural Science Foundation of China (12071273); MOE Project of Humanities and Social Sciences (25YJCZH291); Shanghai Rising-Star Program (22QA1403500); Shanghai Science and Technology Development Funds (23JC1402100); Shanghai Research Center for Data Science and Decision Technology; National Institutes of Health (CA204120 and CA121974); and National Science Foundation (2209685).